A line with a slope of -1 passes through the points (9,-7) and (q,-6). What is the value of q?
q=

thank you

Answer by rothauserc(4718)

slope = (y2 - y1) / (x2 - x1)
:
-1 = (-6 - (-7)) / q - 9)
:
-1 = 1 / (q - 9)
:
-1(q - 9) = 1
:
-q + 9 = 1
:
-q = -8
:
***********
q = 8
**********
: